# Two results on the Convex Algebraic Geometry of sets with continuous symmetries

### v1

@@ 1. Introduction
where $M_{j}\in\operatorname{Sym}^{2}({\mathds{R}}^{d})$ are symmetric $d\times d$ matrices and $M\succeq 0$ means that $M$ is positive-semidefinite. This class of convex sets in ${\mathds{R}}^{n}$ contains all convex polyhedra (as the subclass where all $M_{j}$’s commute), and has received great attention in recent years in the emerging field of Convex Algebraic Geometry, see, e.g. [BPT13, NP23]. Spectrahedra play an important role in Optimization and Applied Mathematics as feasible regions for semidefinite programming, a powerful generalization of linear programming on convex polyhedra to find extrema of linear functions, for which efficient interior-point methods are available. In the literature, spectrahedra with symmetries have been studied both for finite symmetry groups, e.g., [GP04, Val09], and continuous symmetry groups, e.g., [SSS11, KS22, SS20]. In this note, we focus on the continuous case.
@@
The above $\mathsf{G}$-equivariant spectrahedral description for $S$ is built in two steps. First, we define an infinite-dimensional spectrahedral description for $S$, in terms of a $\mathsf{G}$-equivariant linear map $\widetilde{M}\colon{\mathds{R}}^{n}\to\operatorname{Sym}^{2}({\mathcal{V}})$, where ${\mathcal{V}}$ is a Hilbert space, see 3.2. Second, we find a sufficiently large finite-dimensional subspace $V\subset{\mathcal{V}}$ such that $\overline{M}=\widetilde{M}|_{V}$ satisfies $\overline{M}(x)\succeq 0$ if and only if $M(x)\succeq 0$, see 3.5. To demonstrate its practical potential, we carry out this entire construction in three concrete examples, see Examples 3.7, 3.8, and 3.9. For an analogous result concerning finite groups $\mathsf{G}$, see [Kum21].
Our second main result concerns convex sets invariant under a polar representation, i.e., an orthogonal action of a compact Lie group $\mathsf{K}$ on $V$ for which there is a vector subspace $\mathfrak{a}\subset V$ that intersects all $\mathsf{K}$-orbits orthogonally, see 4.1. In this situation, there is a finite group $W\subset\mathsf{O}(\mathfrak{a})$, called the Weyl group, such that the orbit spaces $V/\mathsf{K}$ and $\mathfrak{a}/W$ are isometric. An example is the $\mathsf{O}(n)$-action by conjugation on the space $V=\operatorname{Sym}^{2}({\mathds{R}}^{n})$ of symmetric $n\times n$ matrices. By the Spectral Theorem, all $\mathsf{O}(n)$-orbits intersect the subspace $\mathfrak{a}\subset V$ of diagonal matrices orthogonally, and the action of the symmetric group $W=S_{n}$ permuting entries on $\mathfrak{a}$ is such that $V/\mathsf{K}\cong\mathfrak{a}/W$, see 4.4. By Kostant’s Convexity Theorem (see 4.3 for details), for any polar $\mathsf{K}$-representation $V$, the bijection
[…]
@@
$\mathsf{K}\cdot S$ $\;S.$
maps convex subsets to convex subsets. Thus, it is natural to ask if special classes of invariant convex subsets are preserved by (1.1). While we are unable to confirm our expectation that spectrahedra are mapped to spectrahedra (5.4), we exhibit classes that contain all spectrahedra which are mapped to one another:
@@
For a definition of rigidly convex set, see Section 2. According to the Generalized Lax 2.15, all rigidly convex sets are spectrahedra. While this remains a central open problem in Convex Algebraic Geometry, if true, it would imply by Theorem B (iii) that the bijection (1.1) preserves the class of spectrahedra. Our proof of Theorem B (ii) heavily relies on recent work of Kobert and Scheiderer [KS22]. Earlier work by Sanyal and Saunderson [SS20, Thm. 4.1] established Theorem B (ii) in the special case of the $\mathsf{O}(n)$-action by conjugation on $\operatorname{Sym}^{2}({\mathds{R}}^{n})$.
@@ 2. Preliminaries from Convex Algebraic Geometry
In this section, we recall some basic facts about spectrahedra, spectrahedral shadows, and rigidly convex sets; more comprehensive references on the subject are, e.g., [BPT13, NP23, Vin12]. We denote by $\operatorname{Sym}^{2}({\mathds{R}}^{d})$ the vector space of $d\times d$ real symmetric matrices, and by $\operatorname{I}_{d}\in\operatorname{Sym}^{2}({\mathds{R}}^{d})$ the identity matrix. Given $A,B\in\operatorname{Sym}^{2}({\mathds{R}}^{d})$, we write $A\succeq B$ to indicate that $A-B$ is positive-semidefinite.
@@
Spectrahedra $S$ with nonempty interior $S^{\circ}\neq\emptyset$ are algebraic interiors [Vin12, §2]; in fact, we will recall in 2.14 that an even stronger property holds.
@@
Let $S\subset{\mathds{R}}^{n}$ be an algebraic interior and let $p\in{\mathds{R}}[x]$ be a polynomial of minimal degree such that $S={\mathcal{C}}_{p}(x_{0})$ for some $x_{0}\in{\mathds{R}}^{n}$. Then $p$ is unique (up to multiplication by positive constants) and, for every $q\in{\mathds{R}}[x]$ with $S={\mathcal{C}}_{q}(x^{\prime}_{0})$ for some $x^{\prime}_{0}\in{\mathds{R}}^{n}$, we have $q=p\,h$ for some $h\in{\mathds{R}}[x]$.
The polynomial $p$ in 2.5 is called the defining polynomial of $S$.
@@
Consider the decomposition $p=p_{1}\cdots p_{r}$ of $p$ into irreducible factors. Then, the real zeros of each factor $p_{i}$ lie Zariski-dense in its complex zero set by 2.6. For each $i=1,\ldots,r$ let $m_{i}\in{\mathds{N}}$ be such that $\det(M(x))$ is divisible by $p_{i}(x)^{m_{i}}$ but not by $p_{i}(x)^{m_{i}+1}$. In particular, we can write
@@
We now introduce rigidly convex sets, see [Vin12, §2] for details.
@@
Proposition 2.11 (Theorem 6.3 in [BPT13]).
@@
From 2.11 and the definition of the real zero property, we conclude:
@@
Assume that there is $a\in S_{2}\setminus S_{1}$. Let $b\in S_{1}^{\circ}$ and consider the line segment $[a,b]\subset{\mathds{R}}^{n}$ between $a$ and $b$. Then we have that $[a,b]\cap S_{1}$ is of the form $[c,b]$ for some $c\in(a,b)\cap\partial S_{1}$ because $S_{1}$ is closed and convex. Since $S_{2}$ is a closed convex set with a nonempty interior, every point in its boundary is contained in a proper face [Bar02, Cor. 2.8]. Thus, we have that $c\in\partial S_{1}\cap S_{2}^{\circ}$. However, by 2.11 (i), the polynomial $p$ does not vanish on $c$. ∎
@@
Whether or not the converse of 2.14 holds is an open question:
@@ 3. Equivariant descriptions of invariant spectrahedra
Moreover, any $\mathsf{G}$-invariant affine subspace $S\subset{\mathds{R}}^{n}$ admits a $\mathsf{G}$-equivariant spectrahedral description, since $S=W+v$ where $W$ is a $\mathsf{G}$-invariant linear subspace and $v\in S$ is a fixed vector, e.g., we may take $v$ to be the center of mass of the $\mathsf{G}$-orbit of some $s\in S$. The composition of (3.1) and the translation by $-v$, which is $\mathsf{G}$-equivariant, gives the spectrahedral representation $S=\{x\in{\mathds{R}}^{n}:M(x-v)\succeq 0\}$.
Given a nontrivial $\mathsf{G}$-invariant spectrahedron $S=\{x\in{\mathds{R}}^{n}:M(x)\succeq 0\}$, we shall assume without loss of generality that $0\in{\mathds{R}}^{n}$ is an interior point of $S$ and that $M$ is monic, i.e., $M(0)=\operatorname{I}_{d}$. Indeed, if $S^{\circ}=\emptyset$, then we may replace ${\mathds{R}}^{n}$ with the affine hull $\operatorname{aff}(S)$ of $S$, where the relative interior of $S$ is nonempty. If $S\subset\operatorname{aff}(S)$ admits a $\mathsf{G}$-equivariant spectrahedral description, then so does $S\subset{\mathds{R}}^{n}$ by 3.1. Moreover, we may assume $0\in S^{\circ}$ up to translating by an interior point in $S$ which is fixed by $\mathsf{G}$, e.g., the center of mass of the $\mathsf{G}$-orbit of some $s\in S^{\circ}$. Finally, we may assume that $M(x)$ is monic (cf. 2.2) by replacing it with the $r\times r$ upper left submatrix of $P^{T}M(x)P$, where $P$ is invertible and such that
@@
(3.2) $S_{\mathsf{G}}:=\{x\in{\mathds{R}}^{n}:M(gx)\succeq 0\;\text{ for all }g\in\mathsf{G}\}.$
@@
Set $d=1$ and $n=\binom{m+k}{m}$. We identify ${\mathds{R}}^{n}$ with the vector space ${\mathds{R}}[x_{1},\ldots,x_{m}]_{\leq k}$ of all polynomials of degree at most $k$ in $m$ variables, endowed with the natural action of the orthogonal group $\mathsf{G}=\mathsf{O}(m)$. Consider the monic affine-linear map $M\colon{\mathds{R}}^{n}\to{\mathds{R}}$, given by $M(P)=1-P(1,0,\ldots,0)$, and the halfplane (which is a spectrahedron) $S=\{P\in{\mathds{R}}^{n}:M(P)\leq 0\}$. As a consequence of (3.2) and 3.3, the subset $S_{\mathsf{G}}=\{P\in{\mathds{R}}^{n}:\widetilde{M}(P)\succeq 0\}$ consists of all polynomials $P$ on ${\mathds{R}}^{m}$ of degree at most $k$ that are bounded by $1$ on the unit sphere in ${\mathds{R}}^{m}$. By Scheiderer [Sch18], this set is not a spectrahedral shadow for large enough $m$ and $k$.
Let us describe a similar example in the vector space $\operatorname{Sym}^{2}_{b}(\wedge^{2}{\mathds{R}}^{m})$ of algebraic curvature operators $R\colon\wedge^{2}{\mathds{R}}^{m}\to\wedge^{2}{\mathds{R}}^{m}$, $m\geq 5$, see [BKM21] for details. Set $d=1$ and $n=\frac{1}{12}m^{2}(m^{2}-1)$, and consider the action of $\mathsf{G}=\mathsf{O}(m)$ on $\operatorname{Sym}^{2}_{b}(\wedge^{2}{\mathds{R}}^{m})\cong{\mathds{R}}^{n}$. The map $M(R)=1+\sec_{R}(e_{1}\wedge e_{2})$ is a monic affine-linear map, that gives rise to the spectrahedron $S=\{R\in{\mathds{R}}^{n}:M(R)\geq 0\}$. Since the $\mathsf{G}$-action is transitive on the oriented Grassmannian of $2$-planes in ${\mathds{R}}^{m}$, we have that $S_{\mathsf{G}}=\mathfrak{R}_{\sec\geq-1}(m)$ is the set of curvature operators with sectional curvature bounded below by $-1$, which is not a spectrahedral shadow for $m\geq 5$ by [BKM21, Thm. A].
@@
By 3.3 and the assumption that $S$ is $\mathsf{G}$-invariant, we have that $S=\{x\in{\mathds{R}}^{n}:\widetilde{M}(x)\succeq 0\}$. Let $p$ be the defining polynomial of $S$ as an algebraic interior, and let $\xi\colon{\mathds{R}}^{n}\to{\mathds{R}}^{d}$, $q\in{\mathds{R}}[x]$ and $v\in{\mathds{R}}^{d}$ be as in 2.8. Let $W$ be the span (in the vector space of polynomials of degree less than $d$) of the coordinates of $\xi(g^{-1}x)$ for all $g\in\mathsf{G}$. Choose a basis $F_{1}(x),\ldots,F_{m}(x)$ of $W$ and let
@@
Now we prove that the finite-dimensional linear subspace $V$ of $\mathcal{V}$ that is spanned by the columns of $B(g)$ fulfills the desired property; namely, $\overline{M}:=\widetilde{M}|_{V}$ satisfies $\overline{M}(x)\succeq 0$ if and only if $\widetilde{M}(x)\succeq 0$. Let $\overline{S}:=\{a\in{\mathds{R}}^{n}:\overline{M}(a)\succeq 0\}$; we have to show that $\overline{S}=S$. The inclusion $S\subset\overline{S}$ is clear, since $\widetilde{M}(a)\succeq 0$ implies $\overline{M}(a)\succeq 0$. Therefore, by 2.13 and 2.14, it suffices to prove that the defining polynomial $\overline{p}$ of $\overline{S}$ vanishes on the boundary of $S$. By 2.7, it further suffices to prove that $\overline{p}$ vanishes on $U=\{a\in\partial S:\xi(a)\neq 0\}$. By construction, $F(a)\neq 0$ for all $a\in U$. Let $a\in U$ and consider $\psi\in V$ defined by
@@
The algebraic boundary of $S$ is the zero set of $p(x)=1-x_{1}^{2}-x_{2}^{2}$, and from (2.1), setting $v=(1,0)^{T}$, we obtain that $\xi(x)=(1-x_{1},x_{2})^{T}$ satisfies $M(x)\xi(x)=p(x)v$. Thus, the vector space $W$ from the proof of 3.5 is the $3$-dimensional space of polynomials in ${\mathds{R}}^{2}$ of degree at most $1$. We set $F(x)=(1,-x_{1},-x_{2})^{T}$, and
@@
From the above, one easily sees that $\overline{M}(g(s)x)=g(s)\cdot\overline{M}(x)$ is indeed $\mathsf{G}$-equivariant.
@@
While 3.2 leaves $\overline{M}\colon{\mathds{R}}^{5}\to\operatorname{Sym}^{2}({\mathds{R}}^{3})$ unchanged, it also leads to an $\mathsf{O}(2)$-action on $\operatorname{Sym}^{2}({\mathds{R}}^{3})$ that makes it $\mathsf{O}(2)$-equivariant. Namely, if $g(t)\in\mathsf{O}(2)$ as in (3.8) acts on $Q\in\operatorname{Sym}^{2}({\mathds{R}}^{3})$ by $g(t)\cdot Q=\operatorname{diag}(1,g(2t))\,Q\,\operatorname{diag}(1,g(-2t))$, and acts on $\ell\in{\mathds{R}}^{5}$ by keeping track of the action on the above basis of ${\mathds{R}}[x,y]_{4}$ when it acts by the defining representation on $(x,y)\in{\mathds{R}}^{2}$, then $\overline{M}$ is $\mathsf{O}(2)$-equivariant.
@@
and $\xi=(-a_{12},1+a_{11},-b_{12})^{T}$ satisfies $(M+\operatorname{I}_{3})\xi=p\,v$ where $v=(0,1,0)^{T}$. As in the proof of 3.5, analyzing how $\xi$ behaves under the $\mathsf{G}$-action, we set
@@
and let $V\subset L^{2}(\mathsf{G},{\mathds{R}}^{3})$ be the subspace spanned by the columns of $B(g)$, which form an orthonormal basis with respect to (3.3). Here, the Haar measure on $\mathsf{G}$ is the probability measure $\mathrm{d}\mu(g)=\frac{1}{2\pi^{2}}\mathrm{d}\operatorname{vol}$ proportional to the Riemannian volume form of the unit round metric on $\mathsf{G}\cong S^{3}$. Using (3.6), we compute the Gram matrix of $\overline{M}:=\widetilde{M}|_{V}$ in this basis and, shifting back by $-\operatorname{I}_{4}$, we obtain a $\mathsf{G}$-equivariant spectrahedral description of $S$ as the set of $(a_{11},a_{12},a_{22},b_{12})\in{\mathds{R}}^{4}$ such that

### v2

@@ 1. Introduction
where $M_{j}\in\operatorname{Sym}^{2}({\mathds{R}}^{d})$ are symmetric $d\times d$ matrices and $M\succeq 0$ means that $M$ is positive-semidefinite. This class of convex sets in ${\mathds{R}}^{n}$ contains all convex polyhedra (as the subclass where all $M_{j}$’s commute), and has received great attention in recent years in the emerging field of Convex Algebraic Geometry, see, e.g. [BPT13, NP23]. Spectrahedra play an important role in Optimization and Applied Mathematics as feasible regions for semidefinite programming, a powerful generalization of linear programming on convex polyhedra to find extrema of linear functions, for which efficient interior-point methods are available. In the literature, spectrahedra with symmetries have been studied both for finite symmetry groups, e.g., [GP04, Val09], and continuous symmetry groups, e.g., [SSS11, KS22, SS25]. In this note, we focus on the continuous case.
@@
The above $\mathsf{G}$-equivariant spectrahedral description for $S$ is built in two steps. First, we define an infinite-dimensional spectrahedral description for $S$, in terms of a $\mathsf{G}$-equivariant linear map $\widetilde{M}\colon{\mathds{R}}^{n}\to\operatorname{Sym}^{2}({\mathcal{V}})$, where $\mathcal{V}$} is a Hilbert space, see Construction 3.2. Second, we find a sufficiently large finite-dimensional subspace $V\subset{\mathcal{V}}$ such that $\overline{M}=\widetilde{M}|_{V}$ satisfies $\overline{M}(x)\succeq 0$ if and only if $M(x)\succeq 0$, see Proposition 3.5. To demonstrate its practical potential, we carry out this entire construction in three concrete examples, see Examples 3.7, 3.8, and 3.9. For an analogous result concerning finite groups $\mathsf{G}$, see [Kum21].
Our second main result concerns convex sets invariant under a polar representation, i.e., an orthogonal action of a compact Lie group $\mathsf{K}$ on $V$ for which there is a vector subspace $\mathfrak{a}\subset V$ that intersects all $\mathsf{K}$-orbits orthogonally, see Definition 4.1. In this situation, there is a finite group $W\subset\mathsf{O}(\mathfrak{a})$, called the Weyl group, such that the orbit spaces $V/\mathsf{K}$ and $\mathfrak{a}/W$ are isometric. An example is the $\mathsf{O}(n)$-action by conjugation on the space $V=\operatorname{Sym}^{2}({\mathds{R}}^{n})$ of symmetric $n\times n$ matrices. By the Spectral Theorem, all $\mathsf{O}(n)$-orbits intersect the subspace $\mathfrak{a}\subset V$ of diagonal matrices orthogonally, and the action of the symmetric group $W=S_{n}$ permuting entries on $\mathfrak{a}$ is such that $V/\mathsf{K}\cong\mathfrak{a}/W$, see Example 4.4. By Kostant’s Convexity Theorem (see Proposition 4.3 for details), for any polar $\mathsf{K}$-representation $V$, the bijection
$\big\{\mathsf{K}\text{-invariant subsets of }V\big\}$ $\longleftrightarrow\big\{W\text{-invariant subsets of }\mathfrak{a}\big\}$
@@
$\mathsf{K}\cdot S$ $\,\text{\reflectbox{$\longmapsto$}}\;S.$
maps convex subsets to convex subsets. Thus, it is natural to ask if special classes of invariant convex subsets are preserved by (1.1). While we are unable to confirm our expectation that spectrahedra are mapped to spectrahedra (Conjecture 5.4), we exhibit classes that contain all spectrahedra which are mapped to one another:
@@
For a definition of rigidly convex set, see Section 2. According to the Generalized Lax Conjecture 2.15, all rigidly convex sets are spectrahedra. While this remains a central open problem in Convex Algebraic Geometry, if true, it would imply by Theorem B (iii) that the bijection (1.1) preserves the class of spectrahedra. Our proof of Theorem B (ii) heavily relies on recent work of Kobert and Scheiderer [KS22]. Earlier work by Sanyal and Saunderson [SS25, Thm. 4.1] established Theorem B (ii) in the special case of the $\mathsf{O}(n)$-action by conjugation on $\operatorname{Sym}^{2}({\mathds{R}}^{n})$.
@@ 2. Preliminaries from Convex Algebraic Geometry
In this section, we recall some basic facts about spectrahedra, spectrahedral shadows, and rigidly convex sets; more comprehensive references on the subject are, e.g., [BPT13, NP23, Vin12]. We denote by $\operatorname{Sym}^{2}({\mathds{R}}^{d})$ the vector space of $d\times d$ real symmetric matrices, and by $\operatorname{I}_{d}\in\operatorname{Sym}^{2}({\mathds{R}}^{d})$ the identity matrix. Given $A,B\in\operatorname{Sym}^{2}({\mathds{R}}^{d})$, we write $A\succeq B$ to indicate that ABA-B is positive-semidefinite.
2.1. Spectrahedra and their shadows
Let us recall the definitions of spectrahedron and spectrahedral shadow, and make a few initial observations about them.
@@ 2.1. Spectrahedra and their shadows
Differently from polyhedra, images of spectrahedra under affine-linear maps need not be spectrahedra, which motivates the following definition.
@@
Spectrahedra $S$ with nonempty interior $S^{\circ}\neq\emptyset$ are algebraic interiors [Vin12, §2]; in fact, we will recall in Proposition 2.14 that an even stronger property holds.
We now collect some auxiliary results about algebraic interiors that will be useful in later sections to prove Theorem A in the Introduction.
@@
Let $S\subset{\mathds{R}}^{n}$ be an algebraic interior and let $p\in{\mathds{R}}[x]$ be a polynomial of minimal degree such that $S={\mathcal{C}}_{p}(x_{0})$ for some $x_{0}\in{\mathds{R}}^{n}$. Then $p$ is unique (up to multiplication by positive constants) and, for every $q\in{\mathds{R}}[x]$ with $S={\mathcal{C}}_{q}(x^{\prime}_{0})$ for some $x^{\prime}_{0}\in{\mathds{R}}^{n}$, we have q=phq=p\,h for some $h\in{\mathds{R}}[x]$.
The polynomial $p$ in Lemma 2.5 is called the defining polynomial of $S$.
@@
Finally, the following result is key in the proof of Theorem A in the Introduction, to refine infinite-dimensional equivariant spectrahedral representations into finite-dimensional ones, see Proposition 3.5.
@@
Consider the decomposition $p=p_{1}\cdots p_{r}$ of $p$ into irreducible factors. Then, the real zeros of each factor $p_{i}$ lie Zariski-dense in its complex zero set by Lemma 2.6. For each $i=1,\ldots,r$ let $m_{i}\in{\mathds{N}}$ be such that $\det(M(x))$ is divisible by $p_{i}(x)^{m_{i}}$ but not by $p_{i}(x)^{m_{i}+1}$. In particular, we can write
@@
2.2. Rigidly convex sets
We now introduce real zero polynomials and rigidly convex sets, see [Vin12, §2] for details.
@@ 2.2. Rigidly convex sets
The following foundational result from [HV07] shows that the word “some” in Definition 2.10 can be replaced with “every”, and justifies the alluded convexity.
Proposition 2.11 ([HV07], see Theorem 6.3 in [BPT13]).
@@
From Proposition 2.11 and the definition of the real zero property, we conclude:
@@
We now prove a simple criterion to check if nested rigidly convex sets coincide.
@@
Assume that there is $a\in S_{2}\setminus S_{1}$. Let $b\in S_{1}^{\circ}$ and consider the line segment $a,b]\subset{\mathds{R}}^{n}$ between $a$ and $b$. Then we have that $a,b]\cap S_{1}$ is of the form $[c,b]$ for some $c\in(a,b)\cap\partial S_{1}$ because $S_{1}$ is closed and convex. Since $S_{2}$ is a closed convex set with a nonempty interior, every point in its boundary is contained in a proper face [Bar02, Cor. 2.8]. Thus, we have that $c\in\partial S_{1}\cap S_{2}^{\circ}$. However, by Proposition 2.11 (i), the polynomial $p$ does not vanish on $c$. ∎
The main motivation to study rigidly convex sets comes from the following result.
@@
Whether or not the converse of Proposition 2.14 holds is an open question:
@@ 3. Equivariant descriptions of invariant spectrahedra
Moreover, any $\mathsf{G}$-invariant affine subspace $S\subset{\mathds{R}}^{n}$ admits a $\mathsf{G}$-equivariant spectrahedral description, since $S=W+v$ where $W$ is a $\mathsf{G}$-invariant linear subspace and $v\in S$ is a fixed vector, e.g., we may take $v$ to be the center of mass of the $\mathsf{G}$-orbit of some $s\in S$. The composition of (3.1) and the translation by v-v, which is $\mathsf{G}$-equivariant, gives the spectrahedral representation $S=\{x\in{\mathds{R}}^{n}:M(x-v)\succeq 0\}$.
Given a nontrivial $\mathsf{G}$-invariant spectrahedron $S=\{x\in{\mathds{R}}^{n}:M(x)\succeq 0\}$, we shall assume without loss of generality that $0\in{\mathds{R}}^{n}$ is an interior point of $S$ and that $M$ is monic, i.e., $M(0)=\operatorname{I}_{d}$. Indeed, if $S^{\circ}=\emptyset$, then we may replace ${\mathds{R}}^{n}$ with the affine hull $\operatorname{aff}(S)$ of $S$, where the relative interior of $S$ is nonempty. If $S\subset\operatorname{aff}(S)$ admits a $\mathsf{G}$-equivariant spectrahedral description, then so does $S\subset{\mathds{R}}^{n}$ by Example 3.1. Moreover, we may assume $0\in S^{\circ}$ up to translating by an interior point in $S$ which is fixed by $\mathsf{G}$, e.g., the center of mass of the $\mathsf{G}$-orbit of some $s\in S^{\circ}$. Finally, we may assume that $M(x)$ is monic (cf. Remark 2.2) by replacing it with the $r\times r$ upper left submatrix of $P^{T}M(x)P$, where $P$ is invertible and such that
@@
(3.2) $S_{\mathsf{G}}:=\{x\in{\mathds{R}}^{n}:M(gx)\succeq 0\;\text{ for all }g\in\mathsf{G}\}$
@@
Set $d=1$ and $n=\binom{m+k}{m}$. We identify ${\mathds{R}}^{n}$ with the vector space ${\mathds{R}}[x_{1},\ldots,x_{m}]_{\leq k}$ of all polynomials of degree at most $k$ in $m$ variables, endowed with the natural action of the orthogonal group $\mathsf{G}=\mathsf{O}(m)$. Consider the monic affine-linear map $M\colon{\mathds{R}}^{n}\to{\mathds{R}}$, given by $M(P)=1-P(1,0,\ldots,0)$, and the halfplane (which is a spectrahedron) $S=\{P\in{\mathds{R}}^{n}:M(P)\leq 0\}$. As a consequence of (3.2) and Proposition 3.3, the subset $S_{\mathsf{G}}=\{P\in{\mathds{R}}^{n}:\widetilde{M}(P)\succeq 0\}$ consists of all polynomials $P$ on ${\mathds{R}}^{m}$ of degree at most $k$ that are bounded by $1$ on the unit sphere in ${\mathds{R}}^{m}$. By Scheiderer [Sch18], this set is not a spectrahedral shadow for large enough $m$ and $k$.
Let us describe a similar example in the vector space $\operatorname{Sym}^{2}_{b}(\wedge^{2}{\mathds{R}}^{m})$ of algebraic curvature operators $R\colon\wedge^{2}{\mathds{R}}^{m}\to\wedge^{2}{\mathds{R}}^{m}$, $m\geq 5$, see [BKM21] for details. Set $d=1$ and $n=\frac{1}{12}m^{2}(m^{2}-1)$, and consider the action of $\mathsf{G}=\mathsf{O}(m)$ on $\operatorname{Sym}^{2}_{b}(\wedge^{2}{\mathds{R}}^{m})\cong{\mathds{R}}^{n}$. The map $M(R)=1+\sec_{R}(e_{1}\wedge e_{2})$ is a monic affine-linear map, that gives rise to the spectrahedron $S=\{R\in{\mathds{R}}^{n}:M(R)\geq 0\}$. Since the $\mathsf{G}$-action is transitive on the oriented Grassmannian of $2$-planes in ${\mathds{R}}^{m}$, we have that $S_{\mathsf{G}}=\mathfrak{R}_{\sec\geq-1}(m)$ is the set of curvature operators with sectional curvature bounded below by 1-1, which is not a spectrahedral shadow for $m\geq 5$ by [BKM21, Thm. A].
@@
By Proposition 3.3 and the assumption that $S$ is $\mathsf{G}$-invariant, we have that $S=\{x\in{\mathds{R}}^{n}:\widetilde{M}(x)\succeq 0\}$. Let $p$ be the defining polynomial of $S$ as an algebraic interior, and let $\xi\colon{\mathds{R}}^{n}\to{\mathds{R}}^{d}$, $q\in{\mathds{R}}[x]$ and $v\in{\mathds{R}}^{d}$ be as in Lemma 2.8. Let $W$ be the span (in the vector space of polynomials of degree less than $d$) of the coordinates of $\xi(g^{-1}x)$ for all $g\in\mathsf{G}$. Choose a basis $F_{1}(x),\ldots,F_{m}(x)$ of $W$ and let
@@
Now we prove that the finite-dimensional linear subspace $V$ of $\mathcal{V}$ that is spanned by the columns of $B(g)$ fulfills the desired property; namely, $\overline{M}:=\widetilde{M}|_{V}$ satisfies $\overline{M}(x)\succeq 0$ if and only if $\widetilde{M}(x)\succeq 0$. Let $\overline{S}:=\{a\in{\mathds{R}}^{n}:\overline{M}(a)\succeq 0\}$; we have to show that $\overline{S}=S$. The inclusion $S\subset\overline{S}$ is clear, since $\widetilde{M}(a)\succeq 0$ implies $\overline{M}(a)\succeq 0$. Therefore, by Corollary 2.13 and Proposition 2.14, it suffices to prove that the defining polynomial $\overline{p}$ of $\overline{S}$ vanishes on the boundary of $S$. By Corollary 2.7, it further suffices to prove that $\overline{p}$ vanishes on $U=\{a\in\partial S:\xi(a)\neq 0\}$. By construction, $F(a)\neq 0$ for all $a\in U$. Let $a\in U$ and consider $\psi\in V$ defined by
@@
The algebraic boundary of $S$ is the zero set of $p(x)=1-x_{1}^{2}-x_{2}^{2}$, and from (2.1), setting $v=(1,0)^{T}$, we obtain that $\xi(x)=(1-x_{1},x_{2})^{T}$ satisfies $M(x)\xi(x)=p(x)v$. Thus, the vector space $W$ from the proof of Proposition 3.5 is the $3$-dimensional space of polynomials in ${\mathds{R}}^{2}$ of degree at most $1$. We set $F(x)=(1,-x_{1},-x_{2})^{T}$, and
@@
From the above, one easily sees that $\overline{M}(g(s)x)=g(s)\cdot\overline{M}(x)$, and hence $\overline{M}$ is indeed $\mathsf{G}$-equivariant.
@@
While Construction 3.2 leaves $\overline{M}\colon{\mathds{R}}^{5}\to\operatorname{Sym}^{2}({\mathds{R}}^{3})$ unchanged, it also leads to an $\mathsf{O}(2)$-action on $\operatorname{Sym}^{2}({\mathds{R}}^{3})$ that makes it $\mathsf{O}(2)$-equivariant. Namely, if $g(t)\in\mathsf{O}(2)$ as in (3.8) acts on $Q\in\operatorname{Sym}^{2}({\mathds{R}}^{3})$ by $g(t)\cdot Q=\operatorname{diag}(1,g(2t))\,Q\,\operatorname{diag}(1,g(-2t))$, and acts on $\ell\in{\mathds{R}}^{5}$ by keeping track of the action on the above basis of ${\mathds{R}}[x,y]_{4}$ when it acts by the defining representation on $(x,y)\in{\mathds{R}}^{2}$, then $\overline{M}$ is $\mathsf{O}(2)$-equivariant.
@@
and $\xi=(-a_{12},1+a_{11},-b_{12})^{T}$ satisfies $(M+\operatorname{I}_{3})\xi=p\,v$ where $v=(0,1,0)^{T}$. As in the proof of Proposition 3.5, analyzing how $\xi$ behaves under the $\mathsf{G}$-action, we set
@@
and let $V\subset L^{2}(\mathsf{G},{\mathds{R}}^{3})$ be the subspace spanned by the columns of $B(g)$, which form an orthonormal basis with respect to (3.3). Here, the Haar measure on $\mathsf{G}$ is the probability measure $\mathrm{d}\mu(g)=\frac{1}{2\pi^{2}}\mathrm{d}\operatorname{vol}$ proportional to the Riemannian volume form of the unit round metric on $\mathsf{G}\cong S^{3}$. Using (3.6), we compute the Gram matrix of $\overline{M}:=\widetilde{M}|_{V}$ in this basis and, shifting back by $\operatorname{I}_{4}$, we obtain a $\mathsf{G}$-equivariant spectrahedral description of $S$ as the set of $(a_{11},a_{12},a_{22},b_{12})\in{\mathds{R}}^{4}$ such that